User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair/Presenting: Ellen/…

In Attendance: Jim, Liz, Nick, Lucas, Tiege, Ross, Jamie

Apologies: Jonathan, Arwin, Bryn, Lucia

Absent: Jesse



  • Doodle Poll next meeting time looks like Thursday 1pm (everyone who responded could do this time). Any objections?
    • Moved
  • Anemometer at Yg giving crazy readings (R4752). J-Mac disabled to avoid unnecessary windstows, has this been fixed?
    • No it has not been fixed as of yet
  • Ho - AUA012 experiment did not start, major problems with the set up (Tiege) - something wrong with the proc file. Was this a one off thing, or does it need to be investigated further?
    • Need to talk to schedule people (specific to 26m) - fix catalogue
  • Yg maser unable to be viewed via vnc (R1752). Somehow it fixed itself? Has this happened since/previously?
    • Problem with vnc server, maser kept running fine - maybe due to power glitch/networking issue (3G backup), need to restart vnc server (maybe rig it so it restarts automatically)
  • Noticed that in some experiments, there is uncertainty in the missed scans. This is recorded in the log files that show up when running the end of experiment script so it's very easy to check (if you know the rough time this happened) by the person finishing up.
  • New time? Most of the people have voted now, Thursday 1 pm suits all (
  • Drives on 26m should be fixed now (source=stow is now safe to use) - Lucas has written a lovely wiki page on how to fix it
  • Warren still to organise a module page on wiki
  • Weather station at Yg still doesn't work
    • In the mean time need to wait 30 minutes to update weather info
  • Hb back by next month - maybe usable for T2 next week (ask Brett when setting up)
/home/www/auscope/opswiki/data/pages/operations/agenda2016_08_18.txt · Last modified: 2016/08/18 02:21 by Ellen Manning